Solve and check.
4y − 12 = −3 − 3y

Answer by mananth(16949)

4y-12=-3-3y
add 3y
4y+3y-12=-3-3y+3y
7y-12=-3
add 12
7y-12+12=-3+12
7y=9
divide by7
y=9/7
...
m.ananth@hotmail.ca